# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ Abstract.
Large language models have enabled agents that reason, plan, and interact with tools and environments to accomplish complex tasks. As these agents operate over extended interaction horizons, their effectiveness increasingly depends on adapting behavior to individual users and maintaining continuity across time, giving rise to personalized LLM-powered agents. In such long-term, user-dependent settings, personalization permeates the entire decision pipeline rather than remaining confined to surface-level generation. This survey provides a capability-oriented review of personalized LLM-powered agents. We organize the literature around four interdependent components: profile modeling, memory, planning, and action execution. Using this taxonomy, we synthesize representative methods and analyze how user signals are represented, propagated, and utilized, highlighting cross-component interactions and recurring design trade-offs. We further examine evaluation metrics and benchmarks tailored to personalized agents, summarize application scenarios spanning general assistance to specialized domains, and outline future directions for research and deployment. By offering a structured framework for understanding and designing personalized LLM-powered agents, this survey charts a roadmap toward more user-aligned, adaptive, robust, and deployable agentic systems, accelerating progress from prototype personalization to scalable real-world assistants.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: CSURccs: General and reference Surveys and overviewsccs: Human-centered computingccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Large language models (LLMs) have evolved from passive text generators into general-purpose reasoning systems capable of understanding diverse data, planning actions, and using external tools. Building on these capabilities, LLM-powered agents extend LLMs into integrated systems that decompose complex objectives, invoke tools, interact with dynamic environments, and coordinate with humans or other agents to achieve high-level goals over extended horizons (Acharya et al., 2025; Luo et al., 2025a; Wang et al., 2024c; Tran et al., 2025). This shift marks not only a technical expansion of LLM capabilities but also a conceptual transition toward autonomous, adaptive, and socially grounded intelligent assistants.
@@
As agent systems increasingly support long-term interaction and autonomous decision-making, personalization emerges as a central requirement for maintaining contextual continuity and aligning agent behavior with individual user goals across domains such as education, healthcare, and recommendation (Chu et al., 2025; Wang et al., 2025d; Peng et al., 2025). In LLM-powered agents, personalization extends across the full decision pipeline, shaping how agents interpret user intent, retain information over time, generate plans, and execute actions (Zhang et al., 2024e; Tseng et al., 2024). This system-wide scope introduces new challenges, including handling dynamic and multimodal feedback, maintaining consistency across interacting components, balancing user-specific adaptation with general capabilities, and safeguarding privacy and security. Correspondingly, evaluating personalized agents requires moving beyond static correctness to capture long-term effectiveness, adaptability, and user satisfaction.
@@
Despite growing interest, reviews on personalized LLM-powered agents remain fragmented. Existing studies often concentrate on individual capabilities such as user modeling and memory construction (Wu et al., 2025c; Zhang et al., 2025i), planning and reasoning strategies (Wei et al., 2025), or adaptive interaction mechanisms (Li, 2025; Gao et al., 2025). A unified framework that connects personalization objectives across the full agent lifecycle remains limited. As a result, the interactions among different personalization mechanisms, their system-level trade-offs, and their implications for evaluation and deployment are not yet well understood.
@@
This survey presents a capability-oriented and system-level perspective on personalized LLM-powered agents. Personalization is examined as a distributed property instantiated across four interdependent components: profile modeling, memory management, planning, and action execution. This decomposition enables a systematic analysis of how user-specific signals are represented, propagated, and operationalized throughout the agent lifecycle, from intent understanding to real-world outcomes. As illustrated in Figure 1, when an individual user submits a request, a personalized agent coordinates four interdependent capabilities to generate a tailored response. Profile representations and role definitions shape internal preference modeling, memory schemes organize and retrieve relevant contextual information, planning determines personalized decision strategies, and action execution grounds these decisions in the environment. The resulting interaction outcomes provide feedback that refines internal user representations, forming a closed loop that enables iterative and sustained personalization over time. Building on this taxonomy, we review representative methods, benchmarks, and evaluation protocols, summarize major application domains, and identify open challenges and future research directions. The survey aims to clarify the design space of personalized agents, connect benchmark-driven research with real-world deployment requirements, and provide a structured foundation for developing trustworthy, effective, and scalable personalized agent systems.
@@
    We present an extensive review of recent techniques, benchmarks, and evaluation protocols (§3–§7), highlighting how personalization mechanisms operate within and across agent components and how they are assessed in practice.
@@ 2.2. Personalized LLM-powered Agents
A personalized LLM-powered agent is an LLM-based agent whose internal decision-making is adapted to individual users through user-specific preferences. For each user $u\in\mathcal{U}$, the agent maintains a representation of preferences $p_{u}$, which may encode goals, constraints, habits, or behavioral tendencies inferred from interaction history, feedback, or explicit input.
@@ 2.4. User Preferences
User preferences constitute the core signals that enable personalization in LLM-powered agents by conditioning their internal state, decision-making processes, and user-specific feedback. Prior work commonly categorizes preferences according to their mode of expression (Zhang et al., 2024e; Li et al., 2025c). Explicit preferences are directly specified by users and can be incorporated into the agent without additional inference. Implicit preferences are inferred indirectly from behavioral patterns or contextual cues, reflecting user judgments in a latent or under-specified form.
@@
Figure 3. Example of a two-dimensional taxonomy of user preferences. Preferences are categorized by their expression form (explicit vs. implicit) and semantic function (behavioral vs. topical), illustrated through a multi-turn conversational recommendation scenario.
@@ 2.5. Capability-oriented Taxonomy
    Planning integrates user representations into reasoning processes, shaping decision paths, strategy selection, and prioritization under contextual constraints.
@@
    Action Execution operationalizes personalized decisions through tool invocation and environment interaction, enabling the final outcome.
@@ 3. Profile Modeling
Profile modeling defines the fundamental identity of an agent, including its intrinsic attributes, behavioral tendencies, and operational boundaries (Luo et al., 2025a). In general-purpose agents, this identity is either static or pre-defined externally,established through fixed role instructions and domain expertise (Zhang et al., 2024a; Wu et al., 2024b). In personalized LLM-powered agents, however, profile modeling becomes user-aware and adaptive, incorporating information about the user and dynamically adjusting the agent’s role. To capture this bidirectional relationship, we discuss personalization from two perspectives: user profile modeling and agent role definition.
@@ 3.1. User Profile Modeling
User profile modeling extracts and organizes user-specific signals into a representation that can condition downstream personalization modules. In contrast to generic semantic context, profiles primarily encode personalization cues such as stable traits, long-term preferences, and interaction patterns that determine how an agent should respond or act for a particular user (Wu et al., 2024a). We view a user profile as a structured state constructed from heterogeneous features, where the key design questions are which signals are decision-critical and how they should be aggregated into a usable form. Accordingly, we divide profile modeling into persona-based and response-based approaches, reflecting whether the profile is anchored in user attributes or in user-specific feedback on model outputs.
@@
Persona-based methods infer high-level user attributes that remain informative across contexts. Profiles may be built from explicit self-descriptions and from implicit cues in historical interactions, and are often represented as either a structured attribute vector or a free-form natural-language persona summary. For instance, AlignXpert (Li et al., 2025c) constructs a high-dimensional preference space grounded in psychological and alignment-motivated dimensions, then estimates user-specific positions to form a preference vector. FSPO (Singh et al., 2025) adopts a user-description chain-of-thought mechanism to allocate additional inference effort to preference analysis, producing more fine-grained persona descriptions. Beyond one-shot estimation, several systems iteratively refine latent user attributes through ongoing dialogue. RLPA (Zhao et al., 2025b) and Wang and Yang (2025) update user representations over interaction rounds, while RGMem (Tian et al., 2025) builds evolving profiles by extracting user insights from episodic memory fragments and progressively aggregating them via hierarchical refinement. Relatedly, RLPF (Wu et al., 2025a) optimizes concise, human-readable user summaries using a feedback loop tied to downstream performance, and CURIO (Wan et al., 2025) introduces an intrinsic motivation objective that encourages active inference of latent user types during multi-turn interaction.
@@
Response-based methods model user preferences through how a user evaluates candidate outputs, capturing context-sensitive notions of what constitutes a preferred response. These profiles are typically derived from fine-grained supervision such as pairwise comparisons or implicit behavioral signals, and are designed to generalize preference judgments across prompts and situations. A common approach factorizes individual preference into a shared reward feature basis with user-specific weights, enabling rapid adaptation under sparse user data. For example, RFM (Barreto et al., 2025), PReF (Shenfeld et al., 2025), and LoRe (Bose et al., 2025) represent preferences as linear combinations of shared reward features and infer user-specific coefficients from limited feedback. Complementarily, some methods parameterize user preferences in natural language to improve interpretability and controllability. Janus (Lee et al., 2024) organizes preference dimensions hierarchically from coarse categories to finer-grained value descriptions, while PLUS (Nam et al., 2025) learns text-based preference summaries that condition downstream reward modeling for personalized scoring and generation.
@@ 3.2. Agent Role Definition
Agent role definition determines who the agent is in relation to the user, which can be divided into user-simulated agent definition and adaptive agent definition based on whether the agent is on behalf of the user or not.
User-simulated Agent Definition
In some applications, especially delegation-oriented tasks such as user or dialogue simulation, agents are instantiated to act as simulated users with pre-defined attributes and preferences. These profiles usually function as static role configurations, parameterizing agent behavior for testing or evaluation and can substantially influence task planning and execution (Newsham and Prince, 2025). This one-way profile modeling emulates user traits without learning from real interaction, but remains crucial for creating realistic human–agent scenarios, enabling efficient large-scale experimentation, and ensuring behavioral diversity across synthetic users(Wang et al., 2025b; Li et al., 2025b; Gao et al., 2024b).
Adaptive Agent Definition
In interactive personalized systems, the agent role itself becomes conditional on the user profile. The agent dynamically adjusts its persona, tone, or autonomy based on the user’s characteristics, goals, and trust level, enabling a two-way adaptive process. LD-Agent (Li et al., 2024a) adopts a bidirectional user-agent modeling approach (Xu et al., 2022), utilizing a tunable persona extractor to manage long-term persona bank for both sides. PersonaAgent (Zhang et al., 2025c) optimizes the user-specific system prompt through textual loss feedback between simulated and ground-truth responses, enabling test-time user-preference alignment. PersonaX (Shi et al., 2025c) performs offline clustering and multi-persona profiling, combining with profile retrieval for online decision making. RPP (Mao et al., 2025) proposes an instance-wise prompt optimization framework, which includes both the user historical information and the task guidance for agent, aiming to generate personalized discrete prompts for better individual recommendation performance.
@@ 3.3. Discussion
Role of Profile Modeling
Profile modeling provides the foundational layer of personalization by defining both who the user is through user profiling and who the agent becomes through agent role definition. These two components offer complementary but distinct perspectives: user profiling seeks to construct expressive representations of user traits and preferences, while agent role definition determines how the agent’s operational style should adapt to those user characteristics.
Challenges
Despite their tight conceptual coupling, they face different technical challenges. On the one hand, user profiling is challenged by sparse, noisy, and inconsistent preference signals, as well as the potential conflict between long-term tendencies and context-specific preferences. On the other hand, agent role definition faces the trade-off between flexibility and stability across tasks, avoiding both rigid roles and undesirable role drift. Finally, a broader challenge lies in the weak coupling between profiling and role adaptation. Effective personalization requires bidirectional alignment where the agent must update its role based on refined user profiles, while user profiling should leverage signals arising from the agent’s adaptive behavior.
@@
Existing user profiling methods often yield representations that are static or task-specific, and agent roles that lack controllability or principled adaptation dynamics. Moreover, user profiling and agent role definition are typically developed in isolation, with limited mechanisms for mutual adjustment. Going forward, promising directions include lightweight and interpretable profiling methods that remain robust under sparse signals, more unified frameworks that couple evolving user profiles with stable yet adaptive agent roles.

### v2

@@ Abstract.
Large language models have enabled agentic systems that reason, plan, and interact with tools and environments to accomplish complex tasks. As these agents operate over extended interaction horizons, their effectiveness increasingly depends on adapting behavior to individual users and maintaining continuity across interactions, giving rise to personalized LLM-powered agents (PLAs). In such long-term, user-dependent settings, personalization permeates the entire decision pipeline rather than remaining confined to surface-level response generation. This survey provides a capability-oriented review of personalized LLM-powered agents. Existing work is organized around four interdependent capabilities: profile modeling, memory, planning, and action execution. Using this taxonomy, representative methods are synthesized and analyzed to illustrate how user signals are represented, propagated, and utilized across the agent pipeline, highlighting cross-component interactions and recurring design challenges. Evaluation metrics and benchmarking paradigms tailored to personalized agents are further examined, along with application scenarios ranging from conversational assistants to domain-specific expert systems. By clarifying the design space of personalization in agent systems, this survey provides a structured foundation for developing more user-aligned, adaptive, and deployable LLM-powered agents.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXjournal: CSURccs: General and reference Surveys and overviewsccs: Human-centered computingccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Large language models (LLMs) have evolved from passive text generators into general-purpose reasoning systems capable of understanding diverse data, planning actions, and using external tools. Building on these capabilities, LLM-powered agents extend LLMs into integrated systems that decompose complex objectives, invoke tools, interact with dynamic environments, and coordinate with humans or other agents to achieve high-level goals over extended horizons (Acharya et al., 2025; Luo et al., 2025a; Wang et al., 2024c; Tran et al., 2025). This shift marks not only a technical expansion of LLM capabilities but also a broader shift toward autonomous, adaptive, and socially grounded intelligent assistants.
@@
As agent systems increasingly support long-term interaction and autonomous decision-making, personalization emerges as a central requirement for maintaining contextual continuity and aligning agent behavior with individual user preferences across domains such as education, healthcare, and recommendation (Chu et al., 2025; Wang et al., 2025d; Peng et al., 2025). For personalized LLM-powered agents (PLAs), personalization extends beyond response style and operates across the full decision pipeline, influencing how agents infer user intent, preserve user-relevant information over time, generate personalized plans, and carry out actions in external environments (Zhang et al., 2024e; Tseng et al., 2024). However, this broadened scope also introduces substantial challenges, including integrating dynamic and multimodal feedback, preserving consistency across interacting components, reconciling user-specific adaptation with general competence, and safeguarding privacy and security. As a result, the evaluation of PLAs must move beyond static correctness and instead account for long-term effectiveness, adaptability, and user satisfaction.
@@
Despite the growing interest, reviews on PLAs remain fragmented. Prior surveys often focus on isolated capabilities of general LLM-powered agents or isolated components of personalization, such as user modeling and memory construction (Wu et al., 2025d; Zhang et al., 2025g), planning and reasoning strategies (Wei et al., 2025), or adaptive interaction mechanisms (Li, 2025; Gao et al., 2025). Such capability-specific perspectives have provided valuable insights, but a unified view of how personalization objectives propagate across the full agent lifecycle is still lacking. In particular, the interactions among different personalization mechanisms and their implications for evaluation and deployment remain insufficiently understood.
@@
To address this gap, this survey presents a capability-oriented and system-level perspective on PLAs. We examine personalization as a distributed property instantiated across four interdependent components: profile modeling, memory, planning, and action execution. This decomposition enables a systematic analysis of how user-specific signals are represented, propagated, and operationalized throughout the agent lifecycle, from intent understanding to real-world outcomes. As illustrated in Figure 1, when an individual user submits a request, a PLA coordinates these four capabilities to produce a tailored response. Profile representations and role definitions shape the agent’s internal model of the user, the memory module organizes and retrieves relevant contextual information, planning determines personalized decision strategies, and action execution grounds these decisions in the external environment. The resulting outcomes, together with subsequent user feedback, in turn refine the agent’s internal representation of user preferences, forming a closed loop that enables iterative and sustained personalization over time.
Building on this taxonomy, we review representative methods, benchmarks, and evaluation protocols, summarize major application domains, and identify open challenges and future research directions. The survey aims to clarify the design space of personalized agents, connect benchmark-driven research with real-world deployment requirements, and provide a structured foundation for developing trustworthy, effective, and scalable personalized agent systems.
@@
    We present an extensive review of recent techniques, benchmarks, and evaluation protocols (§3–§7), highlighting how personalization mechanisms operate within and across agent components and how personalization is assessed in practice.
@@ 2.2. Personalized LLM-powered Agents
A personalized LLM-powered agent is an LLM-based agent whose internal pipeline is adapted to individual users through user-specific preferences. For each user $u\in\mathcal{U}$, the agent maintains a representation of the user’s preferences $p_{u}$ inferred from interaction history, feedback, or explicit input.
@@ 2.4. User Preferences
User preferences constitute the core signals that enable personalization in LLM-powered agents by conditioning their internal state, decision-making processes, and user-specific feedback. Prior work commonly categorizes preferences according to their mode of expression (Zhang et al., 2024e; Li et al., 2025d). Explicit preferences are directly specified by users and can be incorporated into the agent without additional inference. Implicit preferences are inferred indirectly from behavioral patterns or contextual cues, reflecting user judgments in a latent or under-specified form.
@@
Figure 3. Example of a two-dimensional taxonomy of user preferences. Preferences are categorized by their expression form (explicit vs. implicit) and semantic function (behavioral vs. topical), illustrated through a multi-turn recommendation scenario.
@@ 2.5. Capability-oriented Taxonomy
Personalization in LLM-powered agents does not arise from a single adaptation module, but from the coordinated operation of multiple capabilities distributed across the agent pipeline. Given the diversity of user data, task settings, and preference types, user-specific signals must be represented, retained, reasoned over, and ultimately realized in external actions and outcomes. We therefore adopt a capability-oriented taxonomy that organizes personalization into four interdependent components. These components are not isolated modules, but complementary functional stages that together support end-to-end personalization in agent systems.
@@
    Planning integrates user-specific information into reasoning processes, shaping decision paths, strategy selection, and prioritization under contextual constraints.
@@
    Action Execution operationalizes personalized decisions through tool invocation and environment interaction, enabling the final personalized outcome.
To further clarify the distinctions among these four capabilities, Table 1 compares them along several shared dimensions, including their typical inputs, temporal scope, and primary objectives, providing a unified analytical lens for the remainder of the survey. In the following sections, we review each capability in detail, while noting that most existing methods focus on only one or a subset of these capabilities rather than the full pipeline.
Table 1. Comparison of the four core capabilities in personalized LLM-powered agents.
Capability Typical Inputs Temporal Scope Primary Objective
Profile Modeling User attributes, behavioral history, and user instructions Lifelong Establish user understanding and support agent-user alignment
Memory Interaction history, user events, contextual records Turn to lifelong Maintain temporal continuity and consistent personalization
Planning Profile information, memory, task context, in-task feedback Turn to task Enable personalized reasoning and decision making
Action Execution Plans, tool states, execution feedback Step to task Realize personalized actions and outcomes
@@ 3. Profile Modeling
In general-purpose agent systems, profiles are primarily used to define the agent itself, including its intrinsic attributes, behavioral tendencies, and operational boundaries (Luo et al., 2025a). Such identities are often static or externally specified through fixed role instructions, domain expertise, or pre-defined functional assignments (Zhang et al., 2024a; Wu et al., 2024b). In PLAs, however, profile modeling becomes explicitly user-centered, modeling the user and dynamically shaping how the agent positions itself when carrying out tasks for that user. In this sense, profile modeling serves as the foundational layer of PLAs, linking user understanding with agent role configuration. Accordingly, we discuss profile modeling from two complementary perspectives: user profile modeling and agent role definition.
@@ 3.1. User Profile Modeling
User profile modeling extracts and organizes user-specific signals into representations that can guide how an agent should respond or act (Wu et al., 2024a). In this survey, we distinguish two major paradigms according to how user preferences are represented: persona-based approaches, which model the user through relatively holistic traits and preference descriptions, and response-based approaches, which model the user through preferences over candidate responses or model behaviors.
@@
Persona-based methods model the user through relatively stable traits, characteristics, and long-term preference descriptions that remain informative across contexts. Such profiles may be constructed from explicit self-descriptions, behavioral history, or historical interactions, and are often represented as structured attribute vectors or free-form natural-language persona summaries. In downstream use, persona-based profiles are often incorporated as prompts or constraints to guide personalized generation and decision making. For instance, AlignXpert (Li et al., 2025d) constructs a high-dimensional preference space grounded in psychological and alignment-related dimensions, while FSPO (Singh et al., 2025) produces more fine-grained persona descriptions through a user-description chain-of-thought mechanism. Beyond such one-shot profiling, several methods iteratively refine user representations through interaction: RLPA (Zhao et al., 2025b) and Wang and Yang (2025) update user representations across rounds, RGMem (Tian et al., 2025) hierarchically aggregates user insights from episodic memory. Relatedly, RLPF (Wu et al., 2025b) optimizes concise, human-readable user summaries using a feedback loop tied to downstream performance, and CURIO (Wan et al., 2025) introduces an intrinsic motivation objective that encourages active inference of latent user types during multi-turn interaction.
@@
Response-based modeling represents user preferences through how users evaluate candidate outputs, thereby capturing more fine-grained and context-sensitive notions of what constitutes a preferred response. Compared with persona-based approaches, which emphasize holistic user descriptions, response-based approaches focus more directly on user satisfaction with model behavior. In downstream use, these representations are naturally suited to feedback-driven correction and refinement of personalized outputs. A common strategy is to factorize individual preference into a shared reward feature basis with user-specific weights, enabling rapid adaptation under sparse user data. RFM (Barreto et al., 2025), PReF (Shenfeld et al., 2025), and LoRe (Bose et al., 2025) follow this paradigm by inferring user-specific coefficients from limited feedback. Complementarily, some methods represent preferences in natural language to improve interpretability and controllability. Janus (Lee et al., 2024) organizes preference dimensions hierarchically from coarse categories to fine-grained value descriptions, while PLUS (Nam et al., 2025) learns text-based preference summaries that condition downstream reward modeling for personalized scoring and generation.
@@ 3.2. Agent Role Definition
While user profile modeling focuses on representing the user, agent role definition determines how the agent should position itself with respect to that user. We distinguish between user-substituting agent definition, where the agent is configured to act on behalf of the user, and user-adaptive agent definition, where the agent dynamically adjusts its role to better serve a real user during interaction.
User-substituting Agent Definition
User-substituting role definition is common in delegation-oriented and simulation scenarios, where agents are expected either to perform tasks from the user’s perspective or to generate user-like behaviors and dialogue data. In these cases, the agent is configured to act on behalf of the user by instantiating the user’s profile as its role configuration. The resulting role prior then guides planning and execution toward personalized objectives (Newsham and Prince, 2025). This form of role definition does not primarily rely on online adaptation to a real user, but remains important for user proxy execution, realistic human-agent simulation, and scalable generation of diverse synthetic users (Wang et al., 2025b; Li et al., 2025b; Gao et al., 2024b).
User-adaptive Agent Definition
In interactive personalized systems, PLAs adapt themselves to better support the user during collaboration. Here, the agent role becomes conditional on the user profile, allowing the agent to adjust its persona, tone, level of autonomy, or interaction strategy according to the user’s profile. This forms a two-way adaptive process in which user understanding informs role configuration, and the adapted role in turn affects subsequent interaction. Representative methods realize user-adaptive role definition through either joint persona modeling or profile-conditioned prompt adaptation. LD-Agent (Li et al., 2025c) adopts a bidirectional user-agent modeling framework (Xu et al., 2022) with a tunable persona extractor and a long-term persona bank for both users and agents. PersonaAgent (Zhang et al., 2025c), PersonaX (Shi et al., 2025c), and RPP (Mao et al., 2025) instead adapt the agent role through user-conditioned prompts or profiles: PersonaAgent optimizes user-specific system prompts via textual loss feedback, PersonaX combines offline multi-persona profiling with online profile retrieval, and RPP generates personalized discrete prompts from historical user information and task guidance.
@@ 3.3. Discussion
Role and Challenges
Profile modeling provides the foundational layer of personalization by determining both how the user is represented and how the agent is configured in response to that representation. They bridge raw user-related evidence with downstream personalized memory, planning, and action. At the same time, they face several tightly coupled challenges. User profiling must first cope with sparse, noisy, and inconsistent preference signals. Beyond signal quality, a separate challenge is determining which user-related features are genuinely relevant for downstream personalization. In parallel, agent role definition must balance adaptability with stability across tasks, avoiding both rigid role assignments and undesirable role drift. Finally, a broader challenge lies in the requirement of effective coupling and bidirectional alignment between user profiling and agent role adaptation.
@@
Existing user profiling methods often produce representations that are either overly static or narrowly task-specific, while agent roles often lack controllability or principled adaptation dynamics. Moreover, user profiling and agent role definition are typically developed in isolation, with limited mechanisms for mutual adjustment. Promising directions therefore include lightweight and interpretable profiling methods that remain robust under sparse or noisy signals, as well as unified frameworks that jointly model evolving user profiles and stable yet adaptive agent roles.